OpenStudy (anonymous):

solve the system of inequalities by graphing y≥-3x-4 y≤1/3x+1

OpenStudy (anonymous):

y>=-3 x-4 y&&-3 x-4 y<=1/(3 x)+1 y>=-3 x-4 y&&-3 x-4 y<=(3 x+1)/(3 x) x<=1/6 (-5-sqrt(5)), y>=(-9 x^2-3 x-1)/(12 x) you can see the steps to this question by visiting http://www.wolframalpha.com/input/?i=y%E2%89%A5-3x-4++y%E2%89%A41%2F3x%2B1